# GRAF: Graph Retrieval Augmented by Facts for Legal Question Answering

### v1

@@ 4.2 Algorithm Description
Our algorithm runs for every given question along with the set of choices it presents. In other words, the input to our algorithm is the question-answer pair and the KG.
@@
Claim Graph. A multiple-choice question is mainly made up of choices that present different claims with different truth values. To pick the most suitable answer from a set of choices, we first decompose each (Question, Choice) pair into their underlying claims by using a LLM in our case, but more lightweight solutions can be adopted with the adequate resources. We called this procedure Cross Claim Extraction, as we are interested in claims whose entities come from distinct places from either the question and the given choice, or both entities, could come from the choice as we do not consider the question body itself to present any untrue claims given the format of the exams. Questions generally present hypothetical scenarios or premises that do not contradict the law. In this sense, we instruct the LLM to receive two distinct texts and extract the claims in the aforementioned fashion.
@@
KG Extraction. We first obtain the domain-specific KG from Law-RoG by using the law branch provided in the question. Since it is often unfeasible to consider the entire graph for inference, we resort to sampling the graph via a procedure which will retrieve the most relevant nodes and edges from the KG using a heuristic. In our case, we use a bag-of-words (BoW) approach and incorporate the BM25 retrieval algorithm (robertson1976relevance). Before inputting the data directly to this algorithm, we preprocess the words by tokenizing and lemmatizing them using the model provided in the SpaCy package for the Romanian language to achieve better, less noise results. After selecting the top k entities in the KG, we proceed to select their vicinity with a BFS at a capped depth. Moreover, we also limit the maximum number of entities that are retrieved during this stage. In our work, we use a depth of 1, select the top 10 entities, nodes, and edges from the KG, and select no more than 50 distinct entities.
@@
KG Encoding. To encode the KG, we first use a language encoder model to embed the entities in the KG. After embedding, we use an adapted version of Graph Attention (GAT) (velickovic2018graph) for further capturing the topological information for each entity. Since the original GAT model was developed only for graphs with no relation encoding, we adapted this model for our purpose to work on Knowledge Graphs (KGs). First, we capture relational topological information by computing the attention coefficients for relations in which the current entity $\boldsymbol{e^{ij}_{E}}$ is involved between the current node and the adjacent edges:
$\boldsymbol{e^{ij}_{E}}=\sigma_{A}(\boldsymbol{(a^{E})^{T}}[\boldsymbol{W_{N}h^{i}_{N}}||\boldsymbol{W_{E}h^{j}_{E}}])$ (5)
@@
and calculate the attention coefficients for the nodes to capture inter-entity relations between the current node and neighboring nodes:
$\boldsymbol{e^{ij}_{N}}=\sigma_{A}(\boldsymbol{(a^{N})^{T}}[\boldsymbol{W_{N}h^{i}_{N}}||\boldsymbol{W_{N}h^{j}_{N}}])$ (6)
[…]
@@
We obtain nodes and edges representations by aggregating the information from the adjacent nodes for each node into $\boldsymbol{h^{{}^{\prime}}_{N}}$ and the information from the adjacent edges for each node into $\boldsymbol{h^{{}^{\prime}}_{E}}$:
@@
Final Score. After encoding the graphs, we will select the relevant information from the provided knowledge given the encoded claims. For this, we compute a relevance matrix by calculating the alignment between each encoded claim from the previous discussion, which we denote $\boldsymbol{h^{i}_{c}}$ and all the encoded law sub-KG relations $\boldsymbol{h^{j}}$. We calculate the alignment using the cosine similarity in the following way:
@@
We separately encode the (Question, Choice) pair and decide what information is better suited in the final decision for the current choice. We employ a self-attention mechanism for this task and provide a score based on the gathered information and the given choice:
@@
The term $\bar{c}$ is the encoding of the current (Question, Choice) pair. Finally, we compute the score:
@@
where $\sigma$ is the sigmoid logistic activation used to provide a probability score.
5 Baselines
In this section, we will describe each baseline model used in our experiments. For encoder models, we opt for similar approaches to ones in IR and RAG (wang2024utilizing). In Appendix LABEL:sec:appendix, more experimental details are discussed. As baselines we proposed QBERT, ColQBERT (manotumruksa2020crossbert), ColBERT (khattab2020colbert), Large Language Models (LLMs), LLM with RAG (lewis2020retrieval), LLM fine-tuned with Low Rank Adaptation Method (LoRA) (hu2021lora). More details regarding these baselines can be consulted in the Appendix LABEL:appendix:baselines.
@@ 5 Baselines
Table 3: Results for Promotion exams
6 Results
In this section, we will present the results of our extensive experimentation and discuss the obtained results.
@@ 6.1 Quantitative Assessment
We compare the baselines with our algorithm, where it makes sense to compare the models. In the case of Promotion exams, with questions with a single correct answer, we can compare all the proposed models as performed in Table 5. Encoder-based solutions do not know apriori how many answers to predict, and thus, we resort to comparing Encoder models and LLM models on exams with more than one possible correct answer separately and provide the results as discussed in Section 4.1. The reported scores are the ones the model would receive on the test, no additional points given for wrong answers. There are slight inconsistencies in the results per law branch because the training was performed on the entire dataset at once. We reported these scores for having a more fine-grained view of how strong each model is in every legal area. Nevertheless, our model still surpasses all of the baselines on most of the legal branches, particularly 6 out of 9. Tables 2.1 and LABEL:tab:encbar show the results for the encoder-based models on the Entrance and Bar examinations respectively. Likewise, Table LABEL:tab:llmentr shows the results for the Entrance examinations and Table LABEL:tab:llmbar for the Bar examinations.
@@
CrossQBERT. As proposed in (manotumruksa2020crossbert), we proceed by taking the question and the entire set of possible choices and concatenate all of them together in the same fashion as above and consider the first 3 separator tokens along with a single FFN with a sigmoid activation function which will output three scores for the same question corresponding to each of the answer choices. In this way, we believe in providing BERT with more context for it to be able to have additional information about the neighboring choices to make a better, more informed decision.
@@
ColBERT. Initially an architecture used for IR (khattab2020colbert), we use it for our problem for its underlying philosophy: aligning textual representations. We thus use a model to encode the question and a model to encode the individual choice, and we use the resulting embeddings to perform cosine similarity.
@@
LLM. We use the LLM’s generalization capabilities (zhao2023survey), having decent performances on tasks in no-data settings and no further training. We perform prompt engineering (chen2023unleashing) and extensively experiment with multiple prompts and, in the end, provide the results for the prompts that provided the best performances.
@@
LLM RAG. We make use of Retrieval Augmented Generation (RAG) gao2023retrieval, to provide LLMs with contextual information that would answer the question or guide the model towards the answer. We use the BM25 retriever (robertson1976relevance) along with the SpaCy package for text normalization (tokenization and lemmatization) to extract the top 10 most relevant documents from the corpus. We take the articles from the corpus and chunk them into 50-word documents. We allow consecutive chunks to overlap by 25 words to maintain the context and not potentially disrupt abruptly the flow of the information presented.
@@
LLM LFT. Finally, for our experiments, we fine-tune these LLMs using the LoRA adaptation method (hu2021lora), which was shown experimentally to match the performances of classic full-parameter fine-tuning. This, together with the previous baseline achieve the best results among the baselines. We opt for LoRA method since our computational resources would not allow a full-parameter fine-tuning of all our proposed LLMs.
Appendix D Models

### v2

@@ 4.2 Algorithm Description
Our proposed algorithm is applied for every given question and each of its choice answers. Concretely, the inputs to GRAF are the question-choice pair and the KG containing entities as nodes and relationships between entities as edges. The algorithm is shown in Figure 1 and Algorithm 1 and will be presented in what follows.
@@
Claim Graph. A multiple-choice question is mainly composed of choices that present different claims with different truth values. We are interested in determining claims whose entities come from (1) either the question and the given choice or (2) both come from the choice. Since, in our setting, questions generally present hypothetical scenarios or premises that do not contradict the law, we do not consider the question body itself to present any untrue claims. Therefore, to pick the most suitable answer, we decompose each (question, choice) pair into their underlying claims using an LLM, similar to Edge et al. (2024); however, any lighter solutions can be adopted with adequate resources. We call this procedure Cross Claim Extraction. An example is presented in Figure 2.
@@
KG Sampling. We obtain the domain-specific KG (in our setting, from Law-RoG) by using the law branch related to the question. Since it is often unfeasible to consider the entire graph for inference due to its size, we resort to sampling the graph via a procedure to retrieve the most relevant nodes and edges using a heuristic. First, we preprocess the words by tokenizing and lemmatizing them using the SpaCy package for the Romanian language to achieve better, less noisy results. Then, we use a bag-of-words (BoW) approach and incorporate the BM25 retriever (Robertson and Jones, 1976) to select the top k entities in the KG. We proceed to select their vicinity with a breadth-first search for a limited depth. We also limit the maximum number of entities retrieved during this stage. In our work, we use a depth of 1, select the top 10 entities, nodes, and edges from the KG, and select no more than 50 distinct entities.
@@
KG Encoding. To encode the KG, we use a language encoder model to embed the entities and the relations resulting in the set of node and edge features $\boldsymbol{h_{N}}$ and $\boldsymbol{h_{E}}$, respectively. Then, we adapt the Graph Attention (GAT) (Veličković et al., 2018) to further capture the topological information for each entity. The original GAT model was developed only for graphs with no relation encoding. Therefore, we transform the set of features considering two different linear transformations parametrized by shared $\boldsymbol{W_{N}}$ and $\boldsymbol{W_{E}}$ for the nodes and edges, respectively. To capture the relational topological information, we compute the attention coefficients for relations $\boldsymbol{e^{ij}_{E}}$ in which the current entity is involved between the current node $i$ and the adjacent edges $j$:
$\boldsymbol{e^{ij}_{E}}=\sigma_{A}(\boldsymbol{(a_{E})^{T}}[\boldsymbol{W_{N}h^{i}_{N}}||\boldsymbol{W_{E}h^{j}_{E}}])$ (1)
@@
and calculate the attention coefficients for the nodes $e^{ij}_{N}$ to capture inter-entity relations between the current node $i$ and neighboring nodes $j$:
$\boldsymbol{e^{ij}_{N}}=\sigma_{A}(\boldsymbol{(a_{N})^{T}}[\boldsymbol{W_{N}h^{i}_{N}}||\boldsymbol{W_{N}h^{j}_{N}}])$ (2)
where $\boldsymbol{a_{N}}$ and $\boldsymbol{a_{E}}$ represent two distinct attention vectors for nodes and edges, respectively. We also use a nonlinearity $\sigma_{A}$ as Veličković et al. (2018), which in our case is the LeakyReLU activation function. The $\cdot^{T}$ operator represents transposition, while ||||| | is the concatenation operator.
@@
We obtain final nodes and edges representations by aggregating the information from the adjacent nodes for each node into $\boldsymbol{h^{{}^{\prime}}_{N}}$ and the information from the adjacent edges for each node into $\boldsymbol{h^{{}^{\prime}}_{E}}$:
@@
Final Score. After encoding the graphs, we select the relevant information from the provided knowledge given the encoded claims. For this, we compute a relevance matrix by calculating the alignment between each encoded claim $\boldsymbol{h^{i}_{c}}$ and all the encoded law sub-KG relations $\boldsymbol{h^{j}}$. We calculate the alignment using the cosine similarity:
@@
We separately encode the (question, choice) pair into $\bar{c}$ and decide what information is better suited in the final decision for the current choice. We employ a self-attention mechanism for this task and provide a score based on the gathered information and the given choice:
@@
Finally, we compute the score:
@@
where $\sigma$ is the sigmoid logistic activation used to provide a probability score, and $W_{final}$ is a learnable parameter.
5 Experiments and Results
5.1 Baselines
In this section, we present the baseline models used in our experiments. For encoder models, we opt for similar approaches to ones in information retrieval (IR) and retrieval augmented generation (RAG) (Wang et al., 2024). In Appendix C, more experimental details are discussed. As baselines, we proposed QBERT, ColQBERT (Manotumruksa et al., 2020), ColBERT (Khattab and Zaharia, 2020), Large Language Models (LLMs), LLM with RAG (Lewis et al., 2020), and LLM fine-tuned with Low-Rank Adaptation Method (LoRA) (Hu et al., 2021). More details regarding these baselines can be consulted in Appendix D, and regarding the language models employed, check Appendix E. The results are presented in terms of test accuracy which is further detailed in Appendix F.
Model Civil Penal Proc Civil Proc Penal Administrative Work Family International Commercial Average
@@ 5.1 Baselines
Table 4: Accuracy results for promotion exams.
@@ 5.2 Quantitative Assessment
We compare the baselines with our algorithm, where it makes sense to compare the models. In the case of promotion exams, with questions with a single correct answer, we can compare all the proposed models as performed in Table 5.1. Encoder-based solutions do not know how many answers to predict a priori. Thus, we resort to comparing encoder models and LLM models on exams with more than one possible correct answer separately and provide the results as discussed in §4.1. The reported scores are the ones the model would receive on the test, no additional points given for wrong answers. There are slight inconsistencies in the results per law branch because the training was performed on the entire dataset at once. We reported these scores for having a more fine-grained view of how strong each model is in every legal area. Nevertheless, our model still surpasses the baselines on 6 out of 9 legal branches. Tables 5.2 and 5.2 show the results for the encoder-based models on the entrance and bar examinations, respectively. Likewise, Table 7 shows the results for the entrance examinations and Table 8 for the bar examinations. More granular results are presented in Appendix G.
Model Civil Penal Proc Civil Proc Penal Average
QBERT 41.18 49.10 37.25 39.22 41.69
CrossQBERT 41.18 49.02 43.14 37.25 42.65
ColBERT 45.10 43.14 50.98 41.18 45.10
[…]
Table 5: Accuracy results on encoder models for entrance exams.
Model Civil Penal Proc Civil Proc Penal Average
QBERT 40.32 32.26 40.32 36.29 37.30
CrossQBERT 41.94 38.52 39.52 34.68 38.67
ColBERT 38.71 33.87 36.29 39.52 37.10
[…]
Table 6: Accuracy results on encoder models for bar exams.
Model Civil Penal Proc Civil Proc Penal Average
LLM 39.21 31.37 15.69 27.45 28.43
LLM RAG 43.14 27.45 21.57 27.45 29.90
LLM LFT 37.25 45.10 41.18 41.18 41.18
Table 7: Accuracy results on LLMs for entrance exams.
Model Civil Penal Proc Civil Proc Penal Average
LLM 21.77 25.00 23.39 18.54 22.18
LLM RAG 26.61 29.03 33.87 20.97 27.62
LLM LFT 31.45 32.26 33.06 36.29 33.27
Table 8: Accuracy results on LLMs for bar exams.
Figure 3: Comparative performances of different backbone encoder models (i.e., BERT-based models) of different sizes and pre-trained on different domains.
5.3 Qualitative Assessment
We provide a comparative study for the promotion exams to improve our approach over the baselines shown in Figure 4. In Figure 3, we depict the achieved improvements in our approach when increasing the size of the backbone language encoder model and when changing from the language model trained on general topics to a model focused on our target legal domain. Pre-trained models on the legal domain achieve improved performances over their general domain pre-trained counterparts. As seen in Table 5.1, our model clearly surpasses all the baselines, including larger ones in size, i.e., the LLMs. We notice that the fine-tuned LLMs tend to pick some particular law branches very well but lag behind RAG in other areas. GRAF represents the best of both worlds and achieves good performance where training allows and also maintains competitive performance in tasks where RAG excels in fine-tuning. Our model clearly surpasses the baselines from the same class, the encoder models, in every setting and every exam. Moreover, it benefits from the knowledge provided in the KG, the claim-aware IR, and the network structure of the extracted claims. Our procedure captures key information from the given question and assesses the relevance of the interconnections between the entities present in the question and law as illustrated in Figure 2. A single LLM would have to come up internally with such assessments, and architecturally, it is not necessarily designed in such a fashion. In addition, providing context to LLMs has the potential of awakening the hallucination phenomenon and the model might not correctly identify which claims are indeed true given the related legal texts.
Model Avg. Accuracy
GRAF 55.61
   w/o claim graph 53.14↓2.47
   w/o KG 53.61↓2.00
   w/o claim graph + emb. sum 49.51↓6.10
Table 9: Results for different ablated components of our framework.
Figure 4: Comparative performances of best baselines and our approach on different law branches.
5.4 Ablation Study
In Table 9, we present the effect of removing different parts of our algorithm in an ablation study. We eliminate the claim graph and the KG, and we experiment with collapsing the KG embeddings via summation (denoted emb. sum). In this sense, we experiment with various ablated components and found that missing KG or claim graphs underperform the LLM baselines. Even though the model uses the KG, the performances are poorer because there could be irrelevant information that would play the role of distractors in the decision-making process. We show that our claim-aware information extraction mechanism is indeed effective for improved retrieval capabilities and, consequently, performances. Merely collapsing the extracted information via summation would still not make up for the missing claim-aware IR but would completely damage the model’s performance.
6 Conclusions
We have introduced new resources for the Romanian legal domain, being the first to construct and release a Romanian legal KG. We also proposed a novel approach for KG IR, namely GRAF, which surpassed the baselines proposed for MCQA.
Future Work. We expect solutions that will improve upon our proposed method in every possible aspect. Also, there might be solutions that could potentially explore dataset augmentation via LLMs. Studies could be conducted on target-domain IR, which may include multiple languages and JuRO, CROL and even the KG Law-RoG could represent good foundational resources. We hope that our work will motivate more exploration of under-represented languages with low resources and thus would motivate finding solutions that would work in low-resource settings in any language.
Limitations
We have released a legal MCQA dataset by gathering questions from all available law examinations across the country, which has enough samples for training but might not be sufficient for training on one individual law branch, which is the reason why we opted for training on the whole dataset.
The goal of our work is to improve the resources and provide a methodological approach to addressing legal question-answering. Since such systems are meant to help users understand the law, they are not yet entirely accurate. In fact, the best average score achieved by our approach is only 60%. Therefore, more research is required in this direction since the legal domain is a sensitive topic when considering machine learning systems in assessing the laws. We want to acknowledge that, considering the current state, deploying such systems would require human validation from legal experts to reduce any risk of providing unlawful responses.
Ethical Considerations
We have collected our dataset from diverse official public portals. To protect this dataset from inadequate use, we decided to license the use of this dataset only for research purposes and under no circumstances should it be used in commercial settings. Our work was performed so as not to rely on external human crowd-workers and to not raise any ethical concerns. The data does not contain any sensitive personal information regarding any identifiable real person. All of the hypothetically presented scenarios do not use any person’s name and instead use anonymized abbreviations. Since the data was collected from the public domain and made available by applicable law by the administrative institutions in question, we release these resources under the CC BY-NC-SA 4.0 licensehttps://creativecommons.org/licenses/by-nc-sa/4.0/, allowed by the active European regulationshttps://eur-lex.europa.eu/eli/dir/2019/790/oj.
Appendix A Dataset Analysis
Figure 5: On the left, the number of samples from each law category in the JuRO dataset. On the right, the class equilibrium is depicted via color variations in the heatmap. The heatmap scores are normalized by dividing each value by the maximum of the respective row.
The domain distribution of the JuRO, along with the distribution of the answers, are presented in Figure 5. Because of the examinations’ format, at most two answers are correct. However, in the case of promotion exams, only one answer is correct. The domains of the questions are procedural civil, procedural penal, penal, civil, work, administration, commercial, family, and international.
Table 10 presents the TF-IDF scores Salton et al. (1975) for JuRO dataset, calculated using the following formula:
$\text{score}(t,C)=\frac{f(t,C)}{|\{w|d\in C,w\in d\}|}\log\frac{|C|}{|\{d|d\in C\ and\ t\in d\}|}$
where:
  • the current term for which we compute the score is denoted by $t$;
  • $C$ is the corpus of documents, each document containing multiple words;
  • $f(t,C)$ is the frequency of the term $t$ relative to the corpus $C$.
We notice a high score for the word “penal” over other words, indicating a possible prevalence of penal-related content in the dataset. Moreover, the terms “case”, “term”, “appeal”, “judgement”, “court”, and “request” indicate procedures.
Word Translation TF-IDF Score
cerere request 0.03507
instanţă court 0.03352
caz case 0.02769
lege law 0.02627
penal penal 0.02622
hotărâre decision 0.02614
termen term 0.02602
apel appeal 0.02565
sine self 0.02553
judecată judgement 0.02486
Table 10: TF-IDF scores for the top ten words in the JuRO dataset.
In Table 11, we report the TF-IDF scores for the CROL corpus. Generally, there are words commonly found in articles, but no word indicates a significant bias towards some specific law area.
Word Translation TF-IDF Score
articol article 0.03262
caz case 0.02550
persoană person 0.02510
lege law 0.02471
bun good 0.02429
alineat paragraph 0.02314
prevedea stipulate 0.02260
drept just/correct/law 0.02213
număr number 0.01965
public public 0.01961
Table 11: TF-IDF Scores for the top ten words from the CROL corpus.
In Figure 6, we depict the tokens distribution for both FLAN-T5 (Raffel et al., 2020) and Mistral (Jiang et al., 2023) models using their tokenizers. The distributions behave approximately the same; the difference is that the Mistral tokenizer tends to use more tokens to represent the text than the FLAN-T5 one. Table 12 shows a detailed distribution of questions for each examination.
Appendix B Model Architectures
where $j>k$ and $t^{0}_{i}$ represents the empty sequence. The $\mathcal{Z}$ function maps a given triplet to a sequence that can be processed by the given probabilistic model, known as the model prompt. The prompt has the purpose of facilitating and guiding the model towards a lower on-average negative log-likelihood and consequently, the correct answer. Even though our work also explores sequence-to-sequence models, the ones we chose in particular generate the output in an autoregressive fashion via the decoder module, and thus, our previous discussion still holds.
Encoder Models. These models showed excellent performances on classification tasks despite their relatively smaller size in practice. They feature a good semantic understanding of a given sequence via their pretraining objectives. For instance, BERT (Devlin et al., 2019) featured word-level and sentence-level pretraining, which allowed it to gain a semantic understanding of language. However, they do not exhibit symbol level correlation (Robinson et al., 2023) as LLMs do, and thus, we resort to using their semantic understanding of textual sequences for outputting a number that represents the degree to which a given choice is correct given a question. We consider two learning goals for these models, the binary cross entropy minimization for models outputting probabilities:
$\mathcal{L}_{1}=-(o^{j}_{i}\log(y^{j}_{i})+(1-o^{j}_{i})\log(1-y^{j}_{i}))$ (12)
where $o^{j}_{i}=\mathbbm{1}_{T_{i}}(C^{j}_{i})$ is the ground truth and $y^{j}_{i}$ is the model output probability. We also use the cosine similarity embedding loss to align a given question with the correct answer choice:
$\mathcal{L}_{2}=(1+o^{j}_{i})(1-y^{j}_{i})+(1-o^{j}_{i})y^{j}_{i}$ (13)
During inference, we consider the question along with the set of choices and select the top $|T_{i}|$ scores in the following way:
$Y_{i}^{*}=\text{TopK}(Y_{i},|T_{i}|)$ (14)
Appendix C Experimental Setup
Training was performed on the entire JuRO dataset for every model, and for testing, we considered the checkpoint with the best evaluation results during the training phase. For encoders, we use BERT-based models that were trained for 50 epochs even though in almost all cases, the best model was found around epoch 10; we used a learning rate of $10^{-7}$ and the AdamW Kingma and Ba (2015) optimizer via vanilla PyTorch. All BERT models were fine-tuned on all parameters. LLMs were fine-tuned for 50 epochs using the Trainer API provided by the transformers library using a learning rate of $10^{-7}$, AdamW optimizer, LoRA Hu et al. (2022) alpha of 32, LoRA rank 64, and 2 warm-up steps. All our experiments were performed on a single NVIDIA A100 80GB to which we had limited and restricted access. We report the results on a single run.
Appendix D Baselines
QBERT. We consider the BERT model (Devlin et al., 2019) and construct the input to the LM by appending the given question and the choice in the following way: [CLS] + question + [SEP] + choice + [SEP]. We then use the classification token to attach a fast forward network (FFN) on top with a sigmoid activation function, which will report a score between 0 and 1 for the correctness of the answer choice.
@@ Appendix D Baselines
CrossQBERT. As proposed by Manotumruksa et al. (2020), we proceed by taking the question and the entire set of possible choices and concatenating them together in the same fashion as for QBERT. We consider the first 3 separator tokens and a single FFN with a sigmoid activation function, which outputs three scores for the same question corresponding to each answer choice. In this way, we provide BERT with more context to have additional information about the neighboring choices to make a better, more informed decision.
@@
ColBERT. Initially, an architecture used for IR (Khattab and Zaharia, 2020), we use it for our task for its underlying philosophy: aligning textual representations. We thus use a model to encode the question and a model to encode the individual choice, and we use the resulting embeddings to perform cosine similarity.
@@
LLM. We use the LLM’s generalization capabilities (Zhao et al., 2023), having decent performances on tasks in no-data settings and no further training. We perform prompt engineering (Chen et al., 2023) and extensively experiment with multiple prompts and, in the end, provide the results for the prompts that provided the best performances. For prompts, see Appendix H and the translations in Appendix I.
@@
LLM RAG. We make use of Retrieval Augmented Generation (RAG) Gao et al. (2023) to provide LLMs with contextual information that would answer the question or guide the model towards the answer. We use the BM25 retriever (Robertson and Jones, 1976) along with the SpaCy package for text normalization (tokenization and lemmatization) to extract the top 10 most relevant documents from the corpus. We take the articles from the CROL corpus and chunk them into 50-word documents. We allow consecutive chunks to overlap by 25 words to maintain the context and not potentially abruptly disrupt the flow of the information presented.
@@
LLM LFT. Finally, for our experiments, we fine-tune these LLMs using the LoRA adaptation method (Hu et al., 2021), which was shown experimentally to match the performances of classic full-parameter fine-tuning. This, together with the previous baseline, achieves the best results among the baselines. We opt for the LoRA method since our computational resources would not allow a full-parameter fine-tuning of all our proposed LLMs.
Appendix E Language Models